# Opinion Dynamics for Utility Maximizing Agents: Exploring the Impact of Resource Penalty

### v1

@@ Abstract
We propose a continuous-time nonlinear model of opinion dynamics with utility-maximizing agents connected via a social influence network. A distinguishing feature of the proposed model is the inclusion of an opinion-dependent resource-penalty term in the utilities, which limits the agents from holding opinions of large magnitude. The proposed utility functions also account for how the relative resources within the social group affect both an agent’s stubbornness and social influence. Each agent myopically seeks to maximize its utility by revising its opinion in the gradient ascent direction of its utility function, thus leading to the proposed opinion dynamics. We show that, for any arbitrary social influence network, opinions are ultimately bounded. For networks with weak antagonistic relations, we show that there exists a globally exponentially stable equilibrium using contraction theory. We establish conditions for the existence of consensus equilibrium and analyze the relative dominance of the agents at consensus. We also conduct a game-theoretic analysis of the underlying opinion formation game, including on Nash equilibria and on prices of anarchy in terms of satisfaction ratios. Additionally, we also investigate the oscillatory behavior of opinions in a two-agent scenario. Finally, simulations illustrate our findings.
@@ 1 Introduction
Opinion dynamics deals with the modeling and mathematical analysis of how beliefs and ideas evolve and spread within social groups or networks over time. As collective opinions have far-reaching implications in diverse sectors such as policy, public health, sociology, finance or economics, it is paramount to understand their drivers and consequences. Even for smaller groups, comprehending opinion forming is a necessary first step toward the management of mixed multi-agent groups, their decision making and subsequent dynamic interactions. While many simpler opinion dynamic models have been developed, modeling and analysis of networked rational agents who react to neighbors’ influence remains limited. In this work, we pay particular attention to the setting where the expressed opinions are limited by the (economic, social or time) resources of an agent within a group. We capture this in our model by including a resource penalty term in the utility functions. In the proposed utility functions, stubbornness and social influencing capabilities of an agent also depend on its relative resources within the social group.
@@
The earliest models of opinion dynamics are the French-DeGroot (FD) [1, 2] and Abelson [3] models, which are consensus dynamics. Taylor’s model [4] and the Friedkin-Johnsen (FJ) model [5] expand upon consensus dynamics by introducing stubborn/prejudiced agents to capture the opinion cleavage phenomenon. These models assume that the topology of the social network is constant. In contrast, the Hegselmann-Krause (HK) model [6] considers a variable social network topology, where an agent’s opinion is influenced only by those within a certain confidence bound. The Altafini model [7] explores different opinion behaviors in the presence of antagonistic relations between agents, using signed graphs. Gossip-based models, such as the Deffuant-Weisbuch (DW) model [8], introduce asynchronous interactions, where a random pair of agents update opinions at each time step. For an in-depth summary of the above models, we refer the readers to [9, 10, 11]. These fundamental models serve as the foundation for many others in the literature. Some examples of recent models include the discrete-time versions of the Altafini model with time-varying signed networks [12, 13]. The work [14] proposes the affine boomerang opinion dynamics model with asynchronous opinion updating, incorporating the phenomenon of the boomerang effect into the dynamics. Unlike [7, 12, 13], in this model [14], the opinions do not converge to zero if the network does not satisfy a structural balance property, but rather exhibits bounded fluctuations. Reference [15] proposes the expressed-private-opinion model in which every agent’s private opinion is influenced by the expressed opinions of its neighbors. The concept of social power was first introduced in [1] in order to identify the most influential agent in the social network. The DeGroot-Friedkin (DF) model [16] integrates the opinion formation process with social power evolution through a reflected self-appraisal mechanism. A generalized DF model was proposed in [17] and an extension of the DF model to stubborn agents was proposed in [18]. The modeling of opinion dynamics through a game-theoretic or a utility maximization approach is still in its early stages, with only a limited number of works published thus far. One such model with stubborn agents, presented in [19], uses a best response dynamics framework in which every agent seeks to minimize a cost function. A game theoretic analysis of the FJ-model was carried out in [20]. In a recent study [21], a dynamic influence maximization game is explored, where a set of competing players allocate their fixed resources over certain individuals (who hold opinions about players) to maximize their utilities in the long term. Reference [22] performs a game-theoretical analysis of the asynchronous HK model. The work [23] captures the co-evolution of opinions and actions taken by the agents, where these employ best response dynamics to choose their actions. The work [24] utilizes a continuous-time non-linear opinion dynamics model to tune the mutual cooperative behavior of agents in a repeated game. Within this framework, agents make strategic decisions relying upon rationality and reciprocity. Another approach, detailed in [25], introduces a discrete-time opinion dynamics model with a game-theoretic structure where the agents incur a combination of conformity and manipulation costs based on the opinions. The aim of each agent is to minimize this cost.
@@
Exploration of how agents’ resources impact their opinions and social influencing capabilities is currently an open question in the opinion dynamics literature. Motivated by this, in the current work we adopt a utility maximization and game-theoretic framework to investigate the diverse effects of agents’ resources on the evolution of their opinions. A preliminary version of this work appeared in [26], where we assumed the underlying social influence network to be complete. The current work extends [26] to the case of any general social network topologies and also allows for pairwise antagonistic relationships among the agents. We also investigate several new questions, such as exponential stability of the equilibria, price of anarchy and periodic evolution of opinions.
@@
    We define agent’s utility function to capture the tradeoffs of internal opinion preferences, attachment (or stubbornness) toward its own opinion, conformity or non-conformity and lastly a resource penalty term. Each of the terms in the utility function is affected by the agent’s resources, which is a representation of the agent’s wealth, time etc. We propose a novel opinion dynamics model from the utility functions based upon the assumption that every agent myopically seeks to maximize its own utility. The opinion formation process can be viewed as a one-shot non-cooperative game played by the agents which we refer to as the opinion formation game.
@@
    In the case of weak antagonistic relationships, we show that the game exhibits a unique Nash equilibrium and the agents’ opinions converge to it under the proposed dynamics starting from any arbitrary initial opinion profile. An agent is said to have weak antagonistic relationships if its stubbornness is greater than the collective influence weights of the neighboring agents with whom it has antagonistic relationships. A special case of weak antagonistic relationships is one where antagonistic relationships are absent in the social network.
@@
    We analyze oscillatory and periodic opinion behavior. We show that it is necessary for an agent to be contrarian enough for the two-agent dynamics to have periodic solutions. We also provide sufficient conditions for a Hopf bifurcation to exist for the two-agent dynamics.
@@
Throughout the paper, we use non-bold letters for denoting scalars, bold lowercase letters for denoting vectors, and bold uppercase letters for denoting matrices. The sets of natural numbers, real numbers, non-negative real numbers and positive real numbers are denoted by $\mathbb{N},\mathbb{R}$, $\mathbb{R}_{\geq 0}$ and $\mathbb{R}_{>0}$, respectively. Let $\mathbf{1}\in\mathbb{R}^{n}$ and $\mathbf{0}\in\mathbb{R}^{n}$ denote a vector with all elements equal to one and zero, respectively. For any vector $\mathbf{z}\in\mathbb{R}^{n}$, $\mathbf{z}^{\top}$ denotes its transpose. For any scalar $a\in\mathbb{R}$, $|a|$ denotes its absolute value. For a set $\mathcal{S}$, $\mathcal{S}^{n}$ denotes the Cartesian product of $\mathcal{S}$ with itself $n$ times. The empty set is denoted by $\varnothing$. We denote the difference of any two sets $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ by $\mathcal{S}_{1}\setminus\mathcal{S}_{2}$. Let $\mathsf{G}:=\big{(}\mathcal{V},\mathcal{L}\big{)}$ be a directed graph, where $\mathcal{V}$ is the set of nodes and $\mathcal{L}$ is the set of directed arcs. In a directed graph $\mathsf{G}$, a directed walk from a node $i_{1}\in\mathcal{V}$ to any node $i_{l}\in\mathcal{V}$ is a sequence of nodes $i_{1}\mapsto i_{2}\mapsto\ldots\mapsto i_{l}$ such that $(i_{s},i_{s+1})\in\mathcal{L},\ \forall s\in\{1,2,\ldots,l-1\}$.
@@ 2 Preliminaries
$\mu(\mathbf{A}):=\lim_{h\to 0^{+}}\frac{\|\mathbf{I}_{n}+h\mathbf{A}\|-1}{h}\>\>\in\mathbb{R}.$
@@
(Equilibrium of a strongly contracting system[27]) Suppose $\mathcal{C}\subset\mathbb{R}^{n}$ is convex, closed and positive $\mathbf{f}$-invariant. If $\mathbf{f}:\mathcal{C}\to\mathbb{R}^{n}$ is strongly infinitesimally contracting on $\mathcal{C}$ with rate $\alpha>0$ then, $\mathbf{f}$ has a unique globally exponentially stable equilibrium $\mathbf{z}^{*}\in\mathcal{C}$. normal-$\bullet$
@@ 3 Modeling and Problem Setup
Consider a set $\mathcal{V}:=\{1,\ldots,n\}$ of $n$ agents, with heterogeneous resources, that form opinions on a single topic. We aim to study the evolution of these opinions as these agents interact with each other over a social network. We start by defining utility functions for each agent, that depend on its preferred opinion, the influence exerted by its neighbors and a resource penalty. Then, we derive the proposed opinion dynamics from the utility functions, assuming that each agent myopically seeks to maximize its utility. Thus, the coupling in the utility functions creates the coupling in the opinion dynamics.
@@
We denote the expressed opinion of agent $i\in\mathcal{V}$ at time $t$ on the topic as $z_{i}(t)\in\mathbb{R}$. For brevity, we omit the time argument wherever there is no confusion. The vector $\mathbf{z}:=[z_{1},\cdots,z_{n}]^{\top}\in\mathbb{R}^{n}$ represents the stacked opinions $z_{i}$ of all agents $i\in\mathcal{V}$. We first present the utility function of agent $i$, then describe the various parameters and provide motivation for the chosen structure.
The complete opinion profile $\mathbf{z}\in\mathbb{R}^{n}$ determines the utility for each agent $i\in\mathcal{V}$ as follows
$U_{i}(\mathbf{z},p_{i})=\frac{-w_{i}r_{i}}{2B}\big{(}z_{i}-p_{i}\big{)}^{2}-\frac{1}{2}\bigg{(}\sum_{k\in\mathcal{V}}\frac{a_{ik}r_{k}}{B}\left(z_{k}-z_{i}\right)^{2}\bigg{)}$
$\qquad\qquad-\frac{1}{4r_{i}}z_{i}^{4},$ (1)
where $z_{i}\in\mathbb{R}$ and $p_{i}\in\mathbb{R}$ are the expressed opinion and internal preference on the topic of agent $i\in\mathcal{V}$. The parameter $w_{i}\in\mathbb{R}_{>0}$ is the importance that agent $i\in\mathcal{V}$ attaches to its internal preference on the topic, while $r_{i}\in\mathbb{R}_{>0}$ is the resource available to agent $i\in\mathcal{V}$ and $a_{ik}\in\mathbb{R}$ is the weight of the influence of agent $k\in\mathcal{V}$ on agent $i\in\mathcal{V}$. We make the following standing assumption about the parameters in (1).
@@
    (Parameters’ signs). For each agent $i\in\mathcal{V}$, $p_{i}\in\mathbb{R}$, $w_{i}\in\mathbb{R}_{>0}$, $r_{i}\in\mathbb{R}_{>0}$ and $a_{ik}\in\mathbb{R}$, $\forall k\in\mathcal{V}$. $\bullet$
Note that we allow for heterogeneous agents where each agent can have different parameters in their utility function. Moreover, in this paper, we assume that an agent’s utility is affected by others via a directed social influence graph $\mathsf{G}:=\big{(}\mathcal{V},\mathcal{L},\mathbf{A}\big{)}$. The elements of the adjacency matrix $\mathbf{A}$ are denoted by $a_{ij}\in\mathbb{R}$. If $a_{ik}\neq 0$ then there exists a directed link from node $k\in\mathcal{V}$ to node $i\in\mathcal{V}$ with link weight $a_{ik}$. This denotes that agent $k$’s opinion influences the opinion of agent $i$. The sign of $a_{ik}$ denotes the type of influence relationship and its magnitude denotes the degree of influence. Using this idea, we give the following definition.
@@
(Neighbor sets of an agent). For each agent $i\in\mathcal{V}$, the set $\mathcal{N}_{i}^{e}:=\{k\in\mathcal{V}\setminus\{i\}\>|\>a_{ik}<0\}$ denotes the set of its enemies and the set $\mathcal{N}_{i}^{f}:=\{k\in\mathcal{V}\setminus\{i\}\>|\>a_{ik}>0\}$ denotes the set of its friends. Further,
@@
denotes the set of neighbors of agent $i\in\mathcal{V}$. normal-$\bullet$
Note that the self loop weights $a_{ii}$, for any $i\in\mathcal{V}$, do not affect the utility of the agents. We may assume them to be zero without loss of generality. The self influence of the agents is captured by the preference term in the utility function. In our model, opinions are limited by agents’ resources. Each agent $i\in\mathcal{V}$ possesses static (in time) resource $r_{i}\in\mathbb{R}_{>0}$, which can represent quantities such as wealth, time, social influence or combinations of them. The parameter $r_{i}$ can be more precisely interpreted as the rate at which agent $i\in\mathcal{V}$ “uses” the resources or exerts influence on other agents. The total resources available to the collection of agents is $B:=\sum_{k\in\mathcal{V}}r_{k}$ and hence $(r_{i}/B)\in(0,1)$ represents the resources available to agent $i\in\mathcal{V}$ relative to the entire social group. This could be a measure of agent $i$’s social influence.
@@
(Motivation for the utility function). As each agent $i\in\mathcal{V}$ myopically seeks to maximize its utility function $U_{i}$, the three terms in (1) represent penalties on agent $i$’s expressed opinion $z_{i}$ arising from three factors. The first term penalizes the opinion for deviating from its internal preference $p_{i}$. This penalty is directly proportional to the importance weight $w_{i}$ as well as the relative resources $(r_{i}/B)$ available to agent $i$. The second set of terms penalizes the agent’s opinion for being far away from or close to its neighbors’ opinions depending upon the type of influence relationships, i.e., on the sign of $a_{ik}$’s. The first term can be given a similar interpretation as the second set of terms by interpreting it as penalty to agent $i$ for deviating its opinion from the opinion of a fully stubborn virtual agent whose opinion is always anchored to $p_{i}$. The importance weight $w_{i}$ can then be thought of as the weight of influence of this virtual agent on $i$. The third term (called the resource penalty) restricts agent $i$ from holding opinions of larger magnitude due to the limited resources available to it. In particular, the greater the resources that agent $i$ has, the larger the magnitude of the opinions it can hold. Similarly, we see that the greater the relative resources $(r_{i}/B)$ of agent $i$, the greater is its influence on the opinions and utility function of other agents $k\in\mathcal{V}$. We choose a quartic term for the resource penalty so as to model the effect of limited resources on holding opinions of large magnitude. Such a model is useful in situations where the opinions are about quantities such as prices or the amount of resources an agent is willing to spend on something. Mathematically, a quartic penalty term dominates the other quadratic terms for large enough $z_{i}$ and hence acts as a soft constraint on resource usage. The coefficient of the first term, $w_{i}(r_{i}/B)$, quantifies the level of stubbornness of the agent. Factoring the coefficient of the first term as $w_{i}(r_{i}/B)$ allows us to separate the agent’s importance (or bias) towards its internal preference from the agent’s ability, due to the relative resources available to it, to choose an opinion closer to its preference. normal-$\bullet$
@@
    (Preferences are fixed parameters). For each agent $i\in\mathcal{V}$, $p_{i}\in\mathbb{R}$ is a fixed parameter. $\bullet$
@@
$\begin{split}\dot{z}_{i}=&-\frac{w_{i}r_{i}}{B}[z_{i}-p_{i}]+\bigg{[}\sum_{k\in\mathcal{V}}\frac{a_{ik}r_{k}}{B}[z_{k}-z_{i}]\bigg{]}-\frac{z_{i}^{3}}{r_{i}}.\end{split}$ (2)

### v2

# Opinion Dynamics for Utility Maximizing Agents: Exploring the Impact of the Resource Penalty

@@ Abstract
We propose a continuous-time nonlinear model of opinion dynamics with utility-maximizing agents connected via a social influence network. A distinguishing feature of the proposed model is the inclusion of an opinion-dependent resource-penalty term in the utilities, which limits the agents from holding opinions of large magnitude. This model is applicable in scenarios where the opinions pertain to the usage of resources, such as money, time, computational resources etc. Each agent myopically seeks to maximize its utility by revising its opinion in the gradient ascent direction of its utility function, thus leading to the proposed opinion dynamics. We show that, for any arbitrary social influence network, opinions are ultimately bounded. For networks with weak antagonistic relations, we show that there exists a globally exponentially stable equilibrium using contraction theory. We establish conditions for the existence of consensus equilibrium and analyze the relative dominance of the agents at consensus. We also conduct a game-theoretic analysis of the underlying opinion formation game, including on Nash equilibria and on prices of anarchy in terms of satisfaction ratios. Additionally, we also investigate the oscillatory behavior of opinions in a two-agent scenario. Finally, simulations illustrate our findings.
@@ 1 Introduction
Opinion dynamics deals with the modeling and mathematical analysis of how beliefs and ideas evolve and spread within social groups or networks over time. As collective opinions have far-reaching implications in diverse sectors such as policy, public health, sociology, finance or economics, it is paramount to understand their drivers and consequences. Even for smaller groups, comprehending opinion forming is a necessary first step toward the management of mixed multi-agent groups, their decision making and subsequent dynamic interactions. While many simpler opinion dynamic models have been developed, modeling and analysis of networked rational agents who react to neighbors’ influence remains limited. In this work, we pay particular attention to the setting where the opinions of the agents are related to the usage of resources for a particular task. Each agent has heterogeneous resources available to them which limits their opinions about resource usage. We capture this in our model by including a resource penalty term in the utility functions.
@@
Many developments in the opinion dynamics literature stem from classical models, including the French-DeGroot (FD) model [1, 2], the Abelson model [3], the Taylor model [4], the Friedkin-Johnsen (FJ) model [5], the Hegselmann-Krause (HK) model [6], the Altafini model [7], and the Deffuant-Weisbuch (DW) model [8].
@@
For an in-depth summary of the above models and survey of the literature, we refer the readers to [9, 10, 11, 12]. These fundamental models serve as the foundation for many others in the literature. Some examples of recent models include the discrete-time versions of the Altafini model with time-varying signed networks [13, 14]. The work [15] proposes the affine boomerang opinion dynamics model with asynchronous opinion updating, incorporating the phenomenon of the boomerang effect into the dynamics. Unlike [7, 13, 14], in this model [15], the opinions do not converge to zero if the network does not satisfy a structural balance property, but rather exhibits bounded fluctuations. Reference [16] proposes the expressed-private-opinion model in which every agent’s private opinion is influenced by the expressed opinions of its neighbors. The concept of social power was first introduced in [1] in order to identify the most influential agent in the social network. The DeGroot-Friedkin (DF) model [17] integrates the opinion formation process with social power evolution through a reflected self-appraisal mechanism. A generalized DF model was proposed in [18] and an extension of the DF model to stubborn agents was proposed in [19]. Some works on opinion dynamics [15, 20] also analyze fluctuations and periodic behavior of opinions.
The modeling of opinion dynamics through a game-theoretic or a utility maximization approach is still in its early stages, with only a limited number of works published thus far. Reference [21] showed that cooperative control problems, such as consensus seeking, can be effectively tackled using game-theoretic methods, particularly through potential and weakly acyclic games. The works [22, 23] interpret the FJ-model as best response dynamics within a game-theoretic framework, where each agent aims to minimize a cost function. The work [24] introduced a general framework for social influence grounded in the psychological concept of cognitive dissonance and demonstrated that various opinion dynamics models can be viewed as best response dynamics in a coordination game, where utilities are determined by dissonance functions. In a recent study [25], a dynamic influence maximization game is explored, where a set of competing players allocate their fixed resources over certain individuals (who hold opinions about players) to maximize their utilities in the long term. Reference [26] performs a game-theoretical analysis of the asynchronous HK model. The works [27, 28] capture the co-evolution of opinions and actions taken by the agents under a game-theoretic framework. The work [29] utilizes a continuous-time non-linear opinion dynamics model to tune the mutual cooperative behavior of agents in a repeated game. Within this framework, agents make strategic decisions relying upon rationality and reciprocity. Another approach, detailed in [30], introduces a discrete-time opinion dynamics model with a game-theoretic structure where the agents incur a combination of conformity and manipulation costs based on the opinions. The aim of each agent is to minimize this cost.
@@
Exploration of how agents’ resources impact their opinions and social influencing capabilities is currently an open question in the opinion dynamics literature. Motivated by this, we adopt a utility maximization and game-theoretic framework in the current work to investigate the limiting effects of agents’ resources on their opinions about its usage. A preliminary version of this work appeared in [31], where we assumed the underlying social influence network to be complete. The current work extends [31] to the case of any general social network topologies and also allows for pairwise antagonistic relationships among the agents. We also investigate several new questions, such as exponential stability of the equilibria, price of anarchy and periodic evolution of opinions.
@@
    We define agent’s utility function to capture the tradeoffs of internal opinion preferences, attachment (or stubbornness) toward its own opinion, conformity or non-conformity and lastly a resource penalty term. The resource penalty term in the utility function depends on the agent’s resources, which is a representation of the agent’s wealth, time etc. In this work, we assume that agents are involved in the opinion formation process to make some decisions about the use of their resources for a particular task. In the literature, most models of opinion dynamics are generic and seem to assume that the topic of discussion itself has no effect on the opinions’ evolution. However, this is not realistic in our setting. Thus, we include a resource penalty term in the utility that keeps the opinions bounded. We propose a novel opinion dynamics model from the utility functions based upon the assumption that every agent myopically seeks to maximize its own utility. We refer to the underlying game as the opinion formation game.
@@
    In the case of weak antagonistic relationships, we show that the game exhibits a unique Nash equilibrium and the agents’ opinions converge to it under the proposed dynamics starting from any arbitrary initial opinion profile. A special case of weak antagonistic relationships is one where antagonistic relationships are absent in the social network.
@@
    We analyze oscillatory and periodic opinion behavior. We show that it is necessary for the agents to have sufficiently strong antagonistic relations for the two-agent dynamics to have periodic solutions. We also provide sufficient conditions for a Hopf bifurcation to exist for the two-agent dynamics.
@@
Throughout the paper, we use non-bold letters for denoting scalars, bold lowercase letters for denoting vectors, and bold uppercase letters for denoting matrices. The sets of natural numbers, real numbers, non-negative real numbers and positive real numbers are denoted by $\mathbb{N},$, ≥0 and >0, respectively. Let $\mathbf{1}\in\mathbb{R}^{n}$ and $\mathbf{0}\in\mathbb{R}^{n}$ denote a vector with all elements equal to one and zero, respectively. For any vector $\mathbf{z}\in\mathbb{R}^{n}$, $\mathbf{z}^{\top}$ denotes its transpose. For any scalar $a\in$, $|a|$ denotes its absolute value. For a set $\mathcal{S}$, $\mathcal{S}^{n}$ denotes the Cartesian product of $\mathcal{S}$ with itself $n$ times. The empty set is denoted by $\varnothing$. We denote the difference of any two sets $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ by $\mathcal{S}_{1}\setminus\mathcal{S}_{2}$. Let $\mathsf{G}:=\big{(}\mathcal{V},\mathcal{L}\big{)}$ be a directed graph, where $\mathcal{V}$ is the set of nodes and $\mathcal{L}$ is the set of directed arcs. In a directed graph $\mathsf{G}$, a directed walk from a node $i_{1}\in\mathcal{V}$ to any node $i_{l}\in\mathcal{V}$ is a sequence of nodes $i_{1}\mapsto i_{2}\mapsto\ldots\mapsto i_{l}$ such that $(i_{s},i_{s+1})\in\mathcal{L},\ \forall s\in\{1,2,\ldots,l-1\}$.
@@ 2 Preliminaries
$\mu(\mathbf{A}):=\lim_{h\to 0^{+}}\frac{\|\mathbf{I}_{n}+h\mathbf{A}\|-1}{h}\>\>\in.$
@@
(Equilibrium of a strongly contracting system[32]) Suppose $\mathcal{C}\subset\mathbb{R}^{n}$ is convex, closed and positive $\mathbf{f}$-invariant. If $\mathbf{f}:\mathcal{C}\to\mathbb{R}^{n}$ is strongly infinitesimally contracting on $\mathcal{C}$ with rate $\alpha>0$ then, $\mathbf{f}$ has a unique globally exponentially stable equilibrium $\mathbf{z}^{*}\in\mathcal{C}$ with global Lyapunov functions $V(\mathbf{z})=\|\mathbf{z}-\mathbf{z}^{*}\|$ and $V(\mathbf{z})=\|\mathbf{f}(\mathbf{z})\|$. $\bullet$
@@ 3 Modeling and Problem Setup
Consider a set $\mathcal{V}:=\{1,\ldots,n\}$ of $n$ agents, with heterogeneous resources, that form opinions on a single topic, which is about the quantity of resources to use for a certain purpose. We aim to study the evolution of these opinions as the agents interact with each other over a social network. We start by defining utility functions for each agent. Then, we discuss the utility function and our motivations behind choosing it. Then, we derive the proposed opinion dynamics from the utility functions, assuming that each agent myopically seeks to maximize its utility. Thus, the coupling in the utility functions creates the coupling in the opinion dynamics.
@@
We denote the expressed opinion of agent $i\in\mathcal{V}$ at time $t$ on the topic as $z_{i}(t)\in$. For brevity, we omit the time argument wherever there is no confusion. The vector $\mathbf{z}:=[z_{1},\cdots,z_{n}]^{\top}\in\mathbb{R}^{n}$ represents the stacked opinions $z_{i}$ of all agents $i\in\mathcal{V}$. We first present the utility function of agent $i$, then describe the various parameters and provide motivation for the chosen structure.
The complete opinion profile $\mathbf{z}\in^{n}$ determines the utility for each agent $i\in\mathcal{V}$ as follows
$U_{i}(\mathbf{z},p_{i})=-\frac{w_{i}}{2}\big{(}z_{i}-p_{i}\big{)}^{2}-\sum_{k\in\mathcal{V}}\frac{a_{ik}}{2}\left(z_{k}-z_{i}\right)^{2}-\frac{1}{4r_{i}}z_{i}^{4},$ (1)
where $z_{i}\in$ and $p_{i}\in$ are the expressed opinion and internal preference on the topic of agent $i\in\mathcal{V}$ respectively. The parameter $w_{i}\in_{>0}$ is the importance that agent $i\in\mathcal{V}$ attaches to its internal preference on the topic, while $r_{i}\in_{>0}$ is the resources available to agent $i\in\mathcal{V}$ and $a_{ik}\in$ is the weight of the influence of agent $k\in\mathcal{V}$ on agent $i\in\mathcal{V}$. We call the three terms in the utility function as the preference term, social term and resource penalty, respectively. We make the following standing assumption about the parameters in (1).
@@
    (Parameters’ signs.) For each agent $i\in\mathcal{V}$, $p_{i}\in$, $w_{i}\in_{>0}$, $r_{i}\in_{>0}$ and $a_{ik}\in$, $\forall k\in\mathcal{V}$. $\bullet$
Note that we allow for heterogeneous agents where each agent can have different parameters in their utility function. Moreover, in this paper, we assume that an agent’s utility is affected by others via a directed social influence graph $\mathsf{G}:=\big{(}\mathcal{V},\mathcal{L},\mathbf{A}\big{)}$. The elements of the adjacency matrix $\mathbf{A}$ are denoted by $a_{ij}\in$. If $a_{ik}\neq 0$ then there exists a directed link from node $k\in\mathcal{V}$ to node $i\in\mathcal{V}$ with link weight $a_{ik}$. This denotes that agent $k$’s opinion influences the opinion of agent $i$. The sign of $a_{ik}$ denotes the type of influence relationship and its magnitude denotes the degree of influence. Using this idea, we give the following definition.
@@
(Neighbor set of an agent.) For each agent $i\in\mathcal{V}$, the set $\mathcal{N}_{i}^{e}:=\{k\in\mathcal{V}\setminus\{i\}\>|\>a_{ik}<0\}$ denotes the set of its enemies and the set $\mathcal{N}_{i}^{f}:=\{k\in\mathcal{V}\setminus\{i\}\>|\>a_{ik}>0\}$ denotes the set of its friends. Further,
@@
denotes the set of neighbors of agent $i\in\mathcal{V}$. $\bullet$
Note that the self loop weights $a_{ii}$, for any $i\in\mathcal{V}$, do not affect the utility of the agents. We may assume them to be zero without loss of generality. The self influence of the agents is captured by the preference term in the utility function.
Discussion About the Utility Function
Overall, we seek to explain opinion evolution in social networks from a utility maximization perspective. In the sequel, we derive opinion dynamics as a gradient ascent by the agents of their utility functions, with respect to their own opinions. The three terms in (1) represent penalties on agent $i$’s expressed opinion $z_{i}$ arising from three factors. The preference term penalizes opinion deviation from the internal preference $p_{i}$. This penalty is directly proportional to the importance weight (or stubbornness) $w_{i}$. The social term penalizes the agent’s opinion for being far away from or close to its neighbors’ opinions depending upon the type of influence relationships, i.e., on the sign of $a_{ik}$’s. These two terms in the utility function lead to the first two terms in the opinion dynamics (2), which is the well known Taylor’s model of opinion dynamics [4], when only non-antagonistic relations are allowed. In the presence of antagonistic relations, the second term in (1) is motivated by the boomerang effect [34]. So, our proposed model differs from the literature primarily in the resource penalty in the utility function (1).
The motivation for the resource penalty comes from the observation that if the topic of discussion is about the resource usage for some purpose, then the resource limitations of the agents must also have an effect on the opinion evolution. In our model, the resource penalty in the utility function restricts agent $i$ from holding opinions of larger magnitude. In order to better motivate our model, consider the following example.
Example 3.2
Consider $n$ agents with heterogeneous resource limitations. Let $z_{i}(t)$, the opinion of agent $i\in\mathcal{V}$ at time $t$, be the maximum amount of resources agent $i$ is willing to spend on buying a particular product, if it were to do the buying at time $t$. Agent $i$ has an internal opinion about the value of the product, denoted by $p_{i}$. Further, $w_{i}$ models the confidence of agent $i$ in its internal opinion. In addition, the agents are also influenced by the opinions of others within their social neighborhood. For example, they might look at product reviews or inquire within their social circles, which can influence their opinion about the value of the product. The weights $a_{ik}$’s model the trust or confidence of agent $i$ has in agent $k$’s ability to correctly value the product. Alternatively, these weights could simply model the degree to which agent $i$ seeks to mimic or disagree with agent $k$. However, its spending is ultimately constrained by its resource limitations, which motivates the resource penalty in (1). $\bullet$
The psychological and social effects of resource constraints on decision making of the agents have been studied in various fields such as psychology, economics, transportation etc. We refer the readers to the papers [35, 36] and references therein for more details. For example,  [35] discusses the effects of financial constraints on consumer behaviors from four different perspectives: resource scarcity, choice restriction, social comparison and environmental uncertainity.
@@
(On the Resource Penalty.) The parameter $r_{i}$ denotes the amount of the resources (such as wealth, time etc.) available to agent $i\in\mathcal{V}$. It could be a proxy for the maximum budget an agent has to spend, such as the amount of money available to buy a certain good, the amount of time or other resources available for doing a task etc. In our work, we assume that the amount of resource $r_{i}$ is static in time. This is reasonable to assume in scenarios where the agents’ opinions are about the usage of their resources itself. The agents actually use their resources only after its opinion converges or after a sufficiently long evolution time of its opinion. Thus, we can think of the resource penalty as a soft constraint on their opinions, which are in turn about how much resources they could expend for a good or a task. The greater the resources that agent $i$ has, the larger the magnitude of the opinions it can hold.
In this work, we choose a quartic resource penalty function; however, a more general class of penalty functions is also acceptable. Mathematically, all that is needed for ensuring the boundedness of opinions is to choose a non-negative resource penalty term that dominates the other terms in the utility (1) for large enough $|z_{i}|$. $\bullet$
@@
    (Preferences are fixed parameters.) For each agent $i\in\mathcal{V}$, $p_{i}\in$ is a fixed parameter. $\bullet$
@@
$\begin{split}\dot{z}_{i}=&-w_{i}[z_{i}-p_{i}]+\sum_{k\in\mathcal{V}}a_{ik}[z_{k}-z_{i}]-\frac{z_{i}^{3}}{r_{i}}.\end{split}$ (2)